Let T be a random variable giving the number of spots that appear on top when a die is rolled. List the elements of the sample space S for the number of spots that appear on top when a die is rolled and assign a value to each sample point. 


1
Expert's answer
2021-02-26T00:58:41-0500

The elements of the sample space S for the number of spots that appear on top when the die is rolled is:

"S = \{ 1,2,3,4,5,6\}"


The table for the random variable and its associated value is:

"\begin{matrix}
 t & 1&2&3&4&5&6\\\hline
 P(T=t) |& \frac{1}{6}&\frac{1}{6}&\frac{1}{6}&\frac{1}{6}&\frac{1}{6}&\frac{1}{6}&
\end{matrix}"


So that:

"\sum P(T=t) = 1"